# Testing Occupational Gender Bias in Language Models: Towards Robust Measurement and Zero-Shot Debiasing

### v2

@@ Abstract
Generated texts from large language models (LLMs) have been shown to exhibit a variety of harmful, human-like biases against various demographics. These findings motivate research efforts aiming to understand and measure such effects. Prior works have proposed benchmarks for identifying and techniques for mitigating these stereotypical associations. However, as recent research pointed out, existing benchmarks lack a robust experimental setup, hindering the inference of meaningful conclusions from their evaluation metrics. In this paper, we introduce a list of desiderata for robustly measuring biases in generative language models. Building upon these design principles, we propose a benchmark called OccuGender, with a bias-measuring procedure to investigate occupational gender bias. We then use this benchmark to test several state-of-the-art open-source LLMs, including Llama, Mistral, and their instruction-tuned versions. The results show that these models exhibit substantial occupational gender bias. We further propose prompting techniques to mitigate these biases without requiring fine-tuning. Finally, we validate the effectiveness of our methods through experiments on the same set of models. [^1] Our code and data are at https://github.com/chenyuen0103/gender-bias.
@@ 1 Introduction
Large language models (LLMs) have emerged as powerful tools achieving impressive performance on a variety of tasks (Devlin et al., 2019; Radford et al., 2019; Raffel et al., 2020; Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023; Jiang et al., 2023). Apart from opportunities for potential applications, researchers have identified critical risks associated with the technology (Bender et al., 2021; Bommasani et al., 2021; Weidinger et al., 2021). Specifically, harms caused by human-like biases and stereotypes associated with genders are encoded in LLMs (Sheng et al., 2019; Lucy and Bamman, 2021; Zhao et al., 2019; Wan et al., 2023; Zack et al., 2024).
To address these issues, researchers have proposed a multitude of benchmarks and measurement setups for identifying these harmful associations (Sheng et al., 2019; Gehman et al., 2020; Webster et al., 2020; Kirk et al., 2021; Nadeem et al., 2021; Dhamala et al., 2021) as well as methods for reducing and controlling them (Sheng et al., 2020; Liang et al., 2021; Schick et al., 2021a; Zhao and Chang, 2020; Thakur et al., 2023). While these lines of work provide valuable insights and raise awareness of potential harms caused by biases, several studies point out the shortcomings in existing benchmarks for measuring the biases in generative language models Blodgett et al. (2021); Akyürek et al. (2022); Goldfarb-Tarrant et al. (2023).
Figure 1: The relationships among the prompt template, job (stereotype), and gender (demographic). Both the job and template influence a language model’s gender prediction. If the jobs and templates are related, it becomes hard to separate the direct effect of a job on gender prediction from the effect that goes through the template (the spurious path job – template – gender). We avoid this spurious correlation by selecting jobs and templates independently, thus removing the confounding of templates.
@@
StereoSet Nadeem et al. (2021) Exp.-only
@@
Table 1: Comparison of OccuGender with existing datasets to test gender bias. OccuGender has five desired properties: (1) avoiding potential confounders, (2) using an objective (Obj.) labeling pipeline circumventing the subjective labels from manual annotations, (3) reducing to a smaller prediction space by predicting demographics given stereotypes, instead of vice versa, (4) testing for both explicit (Exp.) and implicit (Imp.) biases, and (5) including non-binary genders. See detailed analysis of each column/desideratum in Section 2.1-2.5.
@@
In this paper, we propose a list of desiderata for bias-measuring methodologies: (1) Prompts and stereotypes should be formed independently to eliminate the confounding effect of prompt template selection. Figure 1 illustrates a causal graph where stereotype (job) and template are formed independently. (2) The labeling of stereotypes should be objective. Previous works relying on crowdsourcing  (Zhao et al., 2018; Rudinger et al., 2018; Nangia et al., 2020; Felkner et al., 2023) introduce subjective human judgment, which can vary widely. (3) Queries in a benchmark should result in a small prediction space for language models. Since there are more variations in the language used to describe stereotypes than in the language used to describe demographics, prompts should be designed so that the models predict demographics given stereotypes. (4) A benchmark should measure both explicit and implicit biases. We refer to explicit biases as stereotypical statements and implicit biases as statements that assume the stereotypes to be true. (5) A benchmark should be demographically inclusive, so tests for gender bias should include non-binary genders.
@@
Following these principles, we propose OccuGender, a framework for assessing occupational gender bias. OccuGender selects jobs that are dominated by a certain gender from the U.S. Bureau of Labor Statistics independent of template formation. Our prompts ask models to predict gender or gender expression, modeling the distribution of demographics given stereotypes. OccuGender also assesses both explicit and implicit biases and measures probabilities of male, female, and non-binary gender predictions. Table 1 compares OccuGender with popular gender bias benchmarks (Nabi and Shpitser, 2018; Rudinger et al., 2018; Nadeem et al., 2021; Felkner et al., 2023; Jha et al., 2023).
@@
We apply OccuGender to quantify the occupational gender bias exhibited by several state-of-the-art open-sourced LLMs: Llama-3-8B (AI@Meta, 2024), Mistral-7B (Jiang et al., 2023), Llama-2-7B (Touvron et al., 2023), and their corresponding instruction-tuned versions. From the experiments, we observe that these models show strong stereotypical associations between gender and stereotypically gendered jobs.
To mitigate such biases, we explore the effect of debiasing prompts of varying abstraction levels in reducing LLMs’ occupational gender bias, drawing on recent works showing how prompting can alter language models’ behavior (Reif et al., 2022; Wei et al., 2022; Kojima et al., 2022; Wang et al., 2024). Our findings suggest that specific prompts can effectively reduce LLMs’s gender bias in occupations and prompts with a low level of abstraction, those specifying the type of biases and when to remove them, are more effective in mitigating gender bias than prompts with a high level of abstraction.
@@
    We propose five desiderata for bias-measuring methods. Then we review popular gender bias benchmarks to assess how well they meet these criteria.
@@
    We introduce OccuGender, a novel framework for assessing occupational gender bias that adheres to all five desiderata.
@@
    We apply OccuGender to test six open-sourced LLMs. The results indicate substantial associations between gender and stereotypical occupations within these models.
  4. 4.
    We develop debiasing prompts at varying levels of abstraction to mitigate the observed biases. Our empirical results demonstrate that these debiasing prompts effectively reduce biases in language models.
2 Desiderata for Bias Measurement
In this section, we discuss the desiderata of bias measurement frameworks. Building upon these desiderata, we proposed OccuGender, a framework for measuring occupational gender bias (Section 3). In Table 1, we compare OccuGender with existing gender bias benchmarks.
2.1 No Confounding in the Prompts
The spurious correlation caused by prompt templates should be minimized when measuring the association between stereotypes and demographics. Consider tasks where the language models predict stereotypes given a demographic (Nadeem et al., 2021; Rudinger et al., 2018). When generating a prompt, one typically chooses a template and a demographic jointly so that a specific template might appear with only one demographic (Nangia et al., 2020; Felkner et al., 2023). As a result, it is unclear whether the language model’s response is influenced by the demographics or by the template itself. A similar argument holds when predicting demographics given stereotypes.
To address template confounding, one should ideally query the language models on all possible (template, stereotype, demographic) combinations. By averaging the results across all templates for each (stereotype, demographic) pair, we can isolate the causal effect of stereotypes on demographic predictions. In causality, this method is known as the backdoor adjustment (Pearl et al., 2000; Peters et al., 2017; Pearl and Mackenzie, 2018).
In reality, however, covering the entire space of prompt templates is infeasible. We thus propose forming stereotypes independent of the templates, and for the templates we use, all stereotypes and demographics are tested. This approach ensures little dependence between stereotypes and templates.
In OccuGender, the occupations (stereotypes) are chosen based on the U.S. Bureau of Labor Statistics, independent of the template formation. See Section 6 for details and Figure 1 for an illustration.
2.2 Objective Labels
The labeling of stereotypical expressions should be objective. In prior datasets, Nadeem et al. (2021) and Nangia et al. (2020) rely on human annotations for their tasks. Zhao et al. (2018) employs a rule-based strategy for gender swapping, supported by annotators for the OntoNotes development set. Similarly, Zhao et al. (2019) validate their sentences through human evaluations. Jha et al. (2023) undertake a culturally inclusive approach, leveraging a globally diverse pool of annotators, while Felkner et al. (2023) adopt a community-in-the-loop annotation pipeline. The approaches above rely on human judgment, which can be subjective. In OccuGender, we determine the stereotypical jobs for males and females using data from the U.S. Bureau of Labor Statistics, bypassing the issue of subjective stereotype labelling.
2.3 Small Prediction Space
2.4 Measuring Explicit and Implicit Biases
The biases expressed by language models can be categorized into two types, explicit and implicit. For explicit bias, the models state the stereotypes, e.g., “girls tend to be softer than boys” (Nadeem et al., 2021). Implicit bias, on the other hand, occurs when the models use associations between stereotypes and demographics when generating texts, without stating the association. For instance, in the sentence “the physician hired the secretary because he was overwhelmed with clients,” an implicitly biased model might associate the pronoun “he” with “doctor”. Both explicit and implicit biases should be measured. In benchmarks proposed by  Nadeem et al. (2021); Nangia et al. (2020); Jha et al. (2023), explicit bias measurements are predominantly featured, while  (Rudinger et al., 2018) and Zhao et al. (2018) assess both explicit and implicit biases. To this end, OccuGender is more similar to  Rudinger et al. (2018) and Zhao et al. (2018) in that we design prompts to test both explicit and implicit bias.
2.5 Inclusion of Demographics
A benchmark should be inclusive with respect to the demographics. As the ultimate goal of studying biases in language models is to promote diversity and inclusion, we argue that datasets used to assess biases should themselves be inclusive. Existing benchmarks in gender bias, however, often overlook non-binary genders. Felkner et al. (2023) and Dev et al. (2021) pioneer the study of biases against the LGBTQ+ community in language models. In the spirit of their work, OccuGender includes non-binary gender as a target of measurement.
3 OccuGender: Measuring Occupational Gender Bias
While the desiderata in Section 2 are generally applicable, we propose a framework to quantify the degree of occupational gender bias exhibited by language models following these design principles.
3.1 Objective Stereotype Labelling
@@ 3.2 Predicting Genders Given Occupations
In practice, given a job, we provide a prompt $x:=(x_{1},..,x_{l})$ instructing a language model to generate text about the person practicing the given job, for instance “I recently met a [JOB]“. Consequently, we measure the prediction probability of expressions indicating each gender. For example, given a set of $n$ continuations $C_{f}:=\{c^{(1)},..,c^{(n)}\}$ indicating “Female”, where each answer $c^{(i)}:=(c^{(i)}_{1},..,c^{(i)}_{m_{i}})$ is a string of $m_{i}$ tokens, we measure the probability of a model associating the given job with the gender “Female” as
@@
3.3 Assess Explicit and Implicit Biases
Our task prompts are listed in Table 2. Prompt 1 is designed to measure explicit bias, whereas the remaining three prompts are intended to measure implicit bias. This is because the first prompt directly asks for one’s gender given the occupation, while the other three ask for a pronoun. Therefore, we look at the results of these setups separately in our evaluation in Section 6.
A $\tilde{P}_{m}$ or $\tilde{P}_{f}$ value close to 1 indicates that the model is biased toward males or females for a certain occupation. The ideal ratios among $\tilde{P}_{g}$ vary by use cases. For instance, if a study aims to assess biases across all gender categories, then an ideal unbiased model should yield high $\tilde{P}_{d}$ with $\tilde{P}_{m}\approx\tilde{P}_{f}$. On the other hand, if only the binary genders are of interest, an ideal unbiased model should yield $\tilde{P}_{m}\approx\tilde{P}_{f}$ regardless of $\tilde{P}_{d}$.
Table 2: Task prompts and their corresponding gender expressions. We extract the probabilities of each gender expression from language models. Prompt 1 assesses explicit biases while prompts 2-4 assess implicit biases.
Prompts

### v3

# Causally Testing Gender Bias in LLMs: A Case Study on Occupational Bias

@@ Abstract
Generated texts from large language models (LLMs) have been shown to exhibit a variety of harmful, human-like biases against various demographics. These findings motivate research efforts aiming to understand and measure such effects. This paper introduces a causal formulation for bias measurement in generative language models. Based on this theoretical foundation, we outline a list of desiderata for designing robust bias benchmarks. We then propose a benchmark called OccuGender, with a bias-measuring procedure to investigate occupational gender bias. We test several state-of-the-art open-source LLMs on OccuGender, including Llama, Mistral, and their instruction-tuned versions. The results show that these models exhibit substantial occupational gender bias. Lastly, we discuss prompting strategies for bias mitigation and an extension of our causal formulation to illustrate the generalizability of our framework. [^1] Our code and data are at https://github.com/chenyuen0103/gender-bias.
@@ 1 Introduction
Large language models (LLMs) have emerged as powerful tools achieving impressive performance on a variety of tasks (Devlin et al., 2019; Radford et al., 2019; Raffel et al., 2020; Brown et al., 2020; Chowdhery et al., 2022; Touvron et al., 2023; Jiang et al., 2023). Apart from opportunities for potential applications, researchers have identified critical risks associated with the technology (Bender et al., 2021; Bommasani et al., 2021; Weidinger et al., 2021). Specifically, LLMs encode harms caused by human-like biases and stereotypes associated with genders, among others (Sheng et al., 2019; Lucy and Bamman, 2021; Zhao et al., 2019; Wan et al., 2023; Zack et al., 2024).
To address these issues, researchers have proposed a multitude of benchmarks and measurement setups for identifying these harmful associations (Sheng et al., 2019; Gehman et al., 2020; Webster et al., 2020; Kirk et al., 2021; Nadeem et al., 2021; Dhamala et al., 2021) as well as methods for reducing and controlling them (Sheng et al., 2020; Liang et al., 2021; Schick et al., 2021a; Zhao and Chang, 2020; Thakur et al., 2023; Jain et al., 2024). While these lines of work provide valuable insights and raise awareness of potential harms caused by biases, several studies point out the shortcomings in existing benchmarks for measuring the biases in generative language models Blodgett et al. (2021); Akyürek et al. (2022); Goldfarb-Tarrant et al. (2023).
@@
StereoSet Nadeem et al. (2021) Mixed Exp.-only
@@
Table 1: Comparison of OccuGender with existing datasets to test gender bias. OccuGender has five desired properties: (1) no template confounding, (2) using an objective (Obj.) labeling pipeline circumventing the subjective labels from manual annotations, (3) reducing to a smaller prediction space by predicting gender given stereotypes, (4) testing for both explicit (Exp.) and implicit (Imp.) biases, and (5) including non-binary genders.
@@
In this paper, we introduce a causal formulation for bias measurement, quantifying the causal impact of stereotypes on gender predictions in language models. Unlike existing benchmarks that only measure correlations, our causal formulation uses do-interventions to account for confounding factors. Building upon this theoretical framework, we outline a list of desiderata for bias-measuring methodologies: (1) Prompts and stereotypes should be formed independently to eliminate the confounding effect of prompt template selection. Figure 1 (bottom right) illustrates a causal graph where stereotypes and templates are formed independently. (2) The labeling of stereotypes should be objective. Previous works relying on crowdsourcing (Zhao et al., 2018; Rudinger et al., 2018; Nangia et al., 2020; Felkner et al., 2023) introduce subjective human judgment, which can vary widely across the annotators. (3) Queries in a benchmark should result in a small prediction space for language models. Since there are more variations in the language used to describe stereotypes than in the language used to describe demographics, prompts should be designed so that the models predict demographics given stereotypes. (4) A benchmark should measure both explicit and implicit biases. We refer to explicit biases as stereotypical statements and implicit biases as statements that assume the stereotypes to be true. (5) A benchmark should be demographically inclusive, so tests for gender bias should include non-binary genders.
@@
Following these principles, we propose OccuGender, a benchmark for assessing occupational gender bias. OccuGender selects jobs that are dominated by a certain gender from the U.S. Bureau of Labor Statistics, independent of template formation. Our prompts ask models to predict a gender or gender expression, modeling the distribution of demographics given stereotypes. OccuGender also assesses both explicit and implicit biases and measures probabilities of male, female, and non-binary gender predictions. Table 1 compares OccuGender with popular gender bias benchmarks (Nabi and Shpitser, 2018; Rudinger et al., 2018; Nadeem et al., 2021; Felkner et al., 2023; Jha et al., 2023).
@@
We apply OccuGender to quantify the occupational gender bias exhibited by ten state-of-the-art open-sourced LLMs: Llama-2-7B (Touvron et al., 2023), Llama-3-8B (AI@Meta, 2024), Mistral-7B (Jiang et al., 2023), Gemma-7B (Team et al., 2024a), Gemma-2-9B (Team et al., 2024b), and their corresponding instruction-tuned versions. From the experiments, we observe that these models show strong stereotypical associations between gender and stereotypically gendered jobs.
Lastly, we discuss ways of mitigating biases in language models, forming an exciting research area in its own right. Although this paper focuses on occupational gender bias, we extend our causal formulation and desiderata to gender bias in education to demonstrate the generalizability of our framework. Note that in OccuGender we do not define the ideal behavior of a language model, such as predicting all genders with probability, allowing the framework to adapt to different contexts.
@@
    We propose a causal formulation and five desiderata for bias-measuring methods. We review popular gender bias benchmarks to assess how well they meet these criteria.
@@
    We introduce OccuGender, a novel framework for assessing occupational gender bias that adheres to all five desiderata and estimates the causal effect of stereotypic jobs on gender predictions.
@@
    We apply OccuGender to test ten open-sourced LLMs. The results indicate substantial associations between gender and stereotypical occupations within these models.
2 A Causal Formulation of Bias Measurement
To fill in the gap of previous work, we introduce a formal formulation of bias measurement using the framework of causal inference (Pearl, 1995; Pearl et al., 2000; Peters et al., 2017; Pearl and Mackenzie, 2018).
2.1 Problem Setup
Consider a dataset $\mathcal{D}$ composed of prompts to measure gender bias in language models. Each prompt has two components: a stereotype $S\in\mathcal{S}$ and a template $T\in\mathcal{T}$. There is often a potential common factor ($C$) affecting both $S$ and $T$, such as prompt designer’s bias or intention when choosing a (template, stereotype) pair. Given a prompt, the language model then generates a gender expression $G\in\mathcal{G}$. We illustrate this idea with two examples in Figure 1, one from Stereotype Nadeem et al. (2021), and the second from OccuGender, and their corresponding causal graphs in Figure 1. The task of predicting stereotypes given gender can be formulated analogously. Our goal is to isolate the effect of $S$ on $G$ while controlling for the confounding effect introduced by $C$.
Stereoset She was confident in [blank] but afraid to face the boys club in the industry. # variables $T$ $=t$ $S$ =$=\{\textit{confident},\textit{afraid}\}$ $G$ $=$ herself/himself/themselves
[…]
Figure 1: Comparison of examples from Stereoset (Nadeem et al., 2021) and our OccuGender. (Top left) Stereoset example, where the combination of the template and stereotypes above only occurred in this specific prompt. (Top right) OccuGender example, which iterates over the Cartesian product of templates and stereotypes (jobs), ensuring independence between $T$ and $S$. . (Bottom Left) The causal graph of Figure 1. Both the job and template influence a language model’s gender prediction. In many existing benchmarks, there are potential confounders, such as prompt designers’ bias, affecting the template-stereotype combinations. If the jobs and templates are related, it becomes hard to separate the direct effect of a job on gender prediction from the effect that goes through the template (the spurious path $S\leftarrow C\rightarrow T\rightarrow G$). (Bottom Right) The causal graph of Figure 1 . We avoid the spurious correlation by selecting stereotypes and templates independently and covering all (stereotype, template) pairs, thus removing the confounding through templates.
2.2 Introducing do-Interventions
To access the causal effect $S$ on $G$, specifically through the path $S\rightarrow G$ in the causal graphs, we need to eliminate the influence of cofounders $C$. For example, if a template is designed for the female gender (e.g., containing “she”), a language model’s gender prediction could be driven more by the template than the stereotype itself. In this case, the prompt designer’s intention to write about females is a confounder, influencing $T$ and $S$.
[…]
2.3 Causal Effects Estimation
When estimating the causal effect of $S$ on $G$, $\mathbb{P}\left(G\mid do\left(S=s\right)\right)$, it is tempting to use
$\mathbb{P}\left(G\mid S=s\right)$ (1)
$=$ $\sum_{t\in\mathcal{T}}\mathbb{P}(T=t)\mathbb{P}\left(G\mid S=s,T=t\right).$
However, this conditional probability captures both causal and spurious associations between $G$ and $S$ and therefore is only valid when the paths between $S$ and $G$ are causal. In other words, $\mathbb{P}\left(G\mid S=s\right)$ is a correct estimate for $\mathbb{P}\left(G\mid do\left(S=s\right)\right)$ only if $S$ and $G$ are not confounded. In Figure 1, due to confounding, the conditional probability provides little insight into whether language models associate the stereotype with females.
2.4 Eliminating Confounding in OccuGender
In OccuGender, we remove confounding effects between $S$ and $G$ by forming stereotypes and templates independently. This breaks the spurious path $S\leftarrow C\rightarrow T\rightarrow G$ (bottom right of  Figure 1).
Independent $(S,T)$ pairs make the conditional probability $\mathbb{P}\left(G\mid S=s\right)$ a valid estimate for the interventional probability $\mathbb{P}\left(G\mid do\left(S=s\right)\right)$, ensuring that the measured associations reflect the true causal effect.
2.5 Mathematical Comparison
In this section, we show how OccuGender allows causal estimation under our causal formulation. We rewrite the interventional probability:
$\mathbb{P}\left(G|do(S=s)\right)$ (2)
$=$ $\sum_{t\in{\mathcal{T}}}\mathbb{P}\left(G|do(S=s),T=t\right)\mathbb{P}\left(T=t|S=s\right)$
$=$ $\sum_{t\in{\mathcal{T}}}\mathbb{P}\left(G|S=s,T=t\right)\mathbb{P}\left(T=t|S=s\right)$
When $T$ and $S$ are dependent, we cannot further simplify the expression. Fortunately, when $T$ and $S$ are independent, as in OccuGender, we have $\mathbb{P}\left(T=t|S=s\right)=\mathbb{P}\left(T=t\right)\;\forall t,s$, and the last line of LABEL:eq:interv_obs can be reduced:
$\sum_{t\in{\mathcal{T}}}\mathbb{P}\left(G|S=s,T=t\right)\mathbb{P}\left(T=t|S=s\right)$ (3)
$=$ $\sum_{t\in{\mathcal{T}}}\mathbb{P}\left(G|S=s,T=t\right)\mathbb{P}\left(T=t\right)$
where $\mathcal{T}$ is the space of all possible templates. However, it is infeasible to iterate through all templates in $\mathcal{T}$, we, therefore, collect a wide variety of templates $\mathcal{\widehat{T}}$ generated by ChatGPT-4o and approximate the causal effect based on them. In other words, we use the approximation:
$\mathbb{P}\left(G|do(S=s)\right)$ (4)
$\approx$ $\frac{1}{|\mathcal{\widehat{T}}|}\sum_{t\in{\mathcal{\widehat{T}}}}\mathbb{P}\left(G|S=s,T=t\right)$
[…]
3 OccuGender Benchmark
Following the causal formulation of bias, we introduce our OccuGender benchmark that is composed to satisfy the following five desiderata: (1) no template confounding, (2) objective stereotype labeling, (3) predicting genders given occupations, (4) measuring both explicit and implicit biases, and (5) inclusive gender labels.
Here, the first criterion is satisfied by the independent formation of prompt templates from the stereotype. For each of the rest of the criteria, we introduce them in detail in each of the following subsections below.
3.1 Objective Stereotype Labeling
@@ 3.2 Predicting Genders Given Occupations
In practice, given a job, we provide a prompt $x$ instructing a language model to generate a pronoun about the person practicing the given job, such as “I have a friend who is a firefighter, [he/she/they]”. Next, we measure the prediction probability of expressions indicating each gender. For example, given a set of $n$ continuations $C_{f}:=\{c^{(1)},..,c^{(n)}\}$ indicating “Female”, where each answer $c^{(i)}:=(c^{(i)}_{1},..,c^{(i)}_{m_{i}})$ is a string of $m_{i}$ tokens, we measure the probability of a model associating the given job with the gender “Female” as
@@
3.2.0.0.1 Advantage of A Small Prediction Space
3.3 Comprehensive Template Set to Evaluate Both Explicit and Implicit Biases
The biases expressed by language models can be categorized into two types, explicit and implicit. For explicit bias, the models state the stereotypes, e.g., “girls tend to be softer than boys” (Nadeem et al., 2021). Implicit bias occurs when the models use associations between stereotypes and demographics when generating texts without stating the association. For instance, in the sentence “the physician hired the secretary because he was overwhelmed with clients,” an implicitly biased model might associate the pronoun “he” with “doctor”. Both explicit and implicit biases should be measured. In benchmarks proposed by  Nadeem et al. (2021); Nangia et al. (2020); Jha et al. (2023), explicit bias measurements are predominantly featured, while  (Rudinger et al., 2018) and Zhao et al. (2018) assess both explicit and implicit biases.
3.4 Inclusive Gender Labels
A benchmark should be inclusive with respect to the demographics. As the ultimate goal of studying biases in language models is to promote diversity and inclusion, we argue that datasets used to assess biases should themselves be inclusive. Existing benchmarks in gender bias, however, often overlook non-binary genders (Hossain et al., 2023). Felkner et al. (2023) and Dev et al. (2021) pioneer the study of biases against the LGBTQ+ community in language models. In the spirit of their work, OccuGender includes non-binary gender as a target of measurement.
We compose a total of 50 prompts, evenly distributed over explicit and implicit prompts. We illustrate two examples in Table 2, and include the entire template set in Appendix C. The diverse set of templates ($\mathcal{\widehat{T}}=50$) helps better approximate the marginalization step in Eq. 4.
Table 2: Task prompt examples and their corresponding gender expressions. See our extensive list of 50 templates in Appendix C.
Prompt Examples